OpenStudy (anonymous):

WILL FAN AND MEDAL FOR ANSWER. 7. If sinθ=4/5, what is the value of cosθ in quadrant 1? A. 9/25 B. 16/25 C. 3/5 D. 5/4

OpenStudy (anonymous):

careful here

OpenStudy (anonymous):

it is not \[\sin^2(4/5)+\cos^2(x)=1\] it is \[(\frac{4}{5})^2+\cos^2(x)=1\]

OpenStudy (anonymous):

easier to draw a right triangle 

OpenStudy (anonymous):

there is a picture of an angle whose sine is \(\frac{4}{5}\) find the missing side via pythagoras, or just remember the mother of all right triangles 3 - 4 - 5 then you can take any trig ratio you like

OpenStudy (anonymous):

Thank you for pointing out my mistake.